What are the angularspeed and radial acceleration and tangential acceleration of a spaceship neglecting in circular turn of radius of 3220km at constant speed of 2870km per hour

Grade:12th pass

1 Answers

Khimraj
5 years ago
angular speed \omega = v/r = 2870/3220 =  41/46 radian/hr
tangential acceleration = v2/r = 205/23 
Hope it clears..........................................................................